FEL Upgrade Project Weekly Brief
November 12-16, 2001

During the past 2 weeks the FEL was used to generate approximately 100 watts in the THz region, a world
record by 6 orders of magnitude; work at 5 different wavelengths; and obtain good data for 3 new user
groups.  All this was accomplished with a klystron replacement and a photocathode cesiation tossed into the
operations.  The machine is quite repeatable even at the 3rd harmonic of 3.15 um.

Last week we had the third semi-annual review of the FEL upgrade project chaired by John Albertine
(consultant to ONR) and John Eric (AFRL).  The following is a preliminary summary of the review
committees’ comments.  A formal report by the committee will be delivered to ONR, AFRL and DOE in a
few weeks.

WBS 4 (Injector):
Door modifications were checked and seemed to work much better, some parts are out for silver plating.  The
electrical wireways for the cesium channels and charge collection ring were submitted to the shop.  The three
getter pump assemblies for the gun chamber were received.

WBS 6 (RF):
Zone 3 - Essentially complete.  Missing Kirk Key lock on HVPS and testing of arc & IR detectors.  Will be
completed after OPS run.

Zone 4 - Essentially complete.  Missing power supply for 5 watt RF amplifiers and system test of position 8
klystron & circulator.  Will be completed after the OPS run.

Master Oscillator - Equipment for zones 3 & 4 is in-house, but can not be installed until after the Fall OPS run.

Quarter Klystrons - The LCW piping was modified to allow easier connection to the klystron.  New ion pump
controllers for the klystrons were received and connected.

Quarter HVPS - The circuit design of the HVPS is making progress.  Capacitors were ordered for the DC
filter section.  Resistors will be specified shortly.

WBS 8 (Instrumentation):
Design for the VME Interface board is complete and the Digital Isolation Module is in the "layout" phase.  We
have received additional Four Channel Timing/Logic Boards and will begin installing parts and testing after the

A barcode tracking system for FEL Removal has been developed.  All major components in the current
machine will be barcoded for tracking purposes prior to their removal from the vault.  A database for all
components is being designed to be implement as new installations take place.

Motion sensor cameras and time-lapse VCRs will be installed to track the upgrade process on video.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o  We plan to inspect the first coil potting in next Monday at Magnet Enterprises International.
Injector Dipoles (DU/DV)
o  DULY is revising small portions of the final assemblies that need modification.
Arc Dipoles (GY, GX, GQ)
o  The detail drawings of the bend and Reverse Bend Dipoles (GX & GQ) are finished except for a first
    adjustment to the pole tip to increase the width of the narrowest part of the good field region.  The
    magnetic model is being exercised to obtain this result.
o  The 180° Dipole (DY) is in its procurement cycle.  Many vendors have signed up on the web site.
Magnet Measurement
o  The parts for the dipole measurement stand are signed off
o  We continued work on the probe that will measure the sextupole and the octupole.
3 inch quad (QX)
o  The sample aluminum version of the core passed inspection.
o  The coils from Magnet Enterprises International were inspected at New England Techni Coil. Many
    were found to have indicated shorts while none of the original inspection documents show this result.
    One representative of good and bad will be air shipped back and be inspected again at MEI on our visit
    Monday to resolve the discrepancy.
Trim Quad (QT)
o  The revision is out on a final bid cycle.
Sextupole (SF)
o  DULY Research is now modeling the sextupole.
Octupole (OT)
o  Work on this magnet is on furlough.
Corrector Dipoles (DB, DJ)
o  Milhous Co. continues making DB and DJ coils.
Beam Line and Vacuum
o  The task order with AES to design the Arc Chambers is being implemented by Dave Waldman.
o  Design work on girders and stands continued at a high level on the quadrupole telescope in front of the
    first arc and the region between the cryomodules.

WBS 10 (Wiggler):
We cleaned up some punch list items from the vacuum chamber fit up exercise and began reinstallation of the
measurement apparatus to take a few runs to clear up some minor questions concerning the measurement sets
taken to date.

WBS 11 (Optics):
Detailing continues on melting of the optical cavity components to the vacuum vessel.

The optical metrology system design progresses.  We have began to procure parts.

The following work was done to support operations:
We provided optics support to Harvey Rutt of Southampton Univ. on harmonic generation and optics for pulse
laser deposition of zinc chalcogenide. We provided assistance to Jacob Grun of the Naval Research Lab.  We
provided support for our ablation experiments.

Last week was devoted to THz studies and operation near 1.6 microns using a new mirror set.  The laser lased
well at 1.6 microns and beam was provided for LOP-01-043--Absorption measurements in GaLaS
glasses--performed by Harry Rutt.  Jim Boyce ran parasitically to try to see doubly diffracted X-rays.  This
turned out to be very difficult due to frequent changes in the system.  The detection of doubly diffracted X-rays
requires many hours of steady electron beam and lasing.  In the end, no clear signal was seen. Visitors from
Brookhaven National Lab and Lawrence Berkeley National Lab gathered excellent data on the spectrum and
power of the THz radiation from two different ports on the machine.  To demonstrate how strong this radiation
is, we lit a match using the THz radiation.  Efforts by Masahiko Tani from Japan to image using the THz were
not successful due to alignment difficulties.  Many thanks go out to the many people who worked heroically on
this experiment to get a signal for Prof. Tani.

This week was devoted to lasing near 1 micron.  Operations were delayed somewhat by the loss of an injector
klystron on Monday and the need to cesiate the cathode on Thursday.  Despite this, Jacob Grun from NRL got
all the data he needed on LOP-01-049, Athermal annealing of silicon.  Harvey Rutt, working on LOP-01-044,
managed to create what, to all appearances, look like high quality zinc chalcogenide films using ultrafast laser
ablation with the frequency doubled FEL radiation at 525 nm.  In parallel with Harvey Rutt's work we worked
on LOP-01-057--ablation rates at short wavelengths.  Once the injector RF was repaired, the machine has
been quite reliable this week, though the quantum efficiency was initially low.

The current wafer in the gun has delivered over 5300 Coulombs.  This is the last week that it will deliver
electrons.  We plan to give it a comfortable retirement.